MARS DESERT RESEARCH STATION

Sol Summary – March 31st

MDRS Daily Summary Report for SOL 6, 31Mar2017
Person filing out Report: Victoria LaBarre
Position: Crew OX and Crew Chemist
Summary Title: Clean up
Mission Status: Concluding experiments and EVAs and cleaning Hab for next crew
Sol Activity Summary: Packed up robot, clock and bike generator, Microbiologist EVA, cleaned entire Hab
Look-Ahead Plan: request ability or help in handling toilet issue/End simulation and go back to Grand Junction.
Anomalies in work:  Nothing out of the ordinary
Weather:
-Morning: Overcast, windy
-Afternoon-Sunny, windy, showers
Crew Physical Status: Crew is healthy

Commander Report – March 31st

Crew Commander Report by Otsmar Villarroel/Crew 177 Commander
March 31st/2017
Today, March 31st is our mission’s last day.  This week has been definitely unique.  Thanks to McLennan Community College, Mars Desert Research Station, and my crewmates for this opportunity.
Today afternoon was used to conduct an EVA led by Pitchaya Jingjit.  She collected twenty-one different soil samples from five different places of interest around the MDRS location with the purpose of bringing back to McLennan Community College, in Waco-Texas in order to find bacteria producing antibiotics to combat antibiotic resistance crisis.  Jingjit was accompanied by Caleb Li, Esteban Ramirez, and Otsmar Villarroel.
Geologists, Espinoza and Li, worked in the science lab during the afternoon.
They use the regolith samples collected during geology EVAs and performed Sieving analysis on each of their samples.  They use different trays, each with a different mesh size.  Back on McLennan Community College campus, they are planning to perform spot test analysis for the following: Carbonate properties, oxide, iron, Uranium, and Potassium.
Dinner consisted of salmon and mashed potatoes.  After dinner crew members started cleaning the MDRS facilities and already started to pack and getting ready to return to Earth with the satisfaction of being part of this experience.
See you later Mars! And one more time, thanks to everybody that makes it an awesome experience.

Geology Report – March 30th

rew 177 Geology Report  March 30, 2017
Crew 177 Geologist Report
Dear Mission Support,
Today is a sunny day, and we are ready to head to the geology stop “The Little Canyon”. After the delicious pancake breakfast, Becky, Esteban, Pitchayaba, and I went on the two Rovers and one ATVs with the measure tools – rope and tape.
Our mission for today’s EVA is to measure the height of the little canyon. About 15 mins, we have arrived the little canyon, and I climbed to the top of the canyon and throw the rope down to the cliff, but I held the one end of the rope, and then Esteban made a mark on the rope. We took some great pictures and videos over there. After we came back to the HAB, we measured the length of the rope, the height of the little canyon is around 18 feet. It was a great experience.
Caleb Li

EVA Report – March 30th

EVA Report:
EVA # 7
Crew members:  Caleb Li, Esteban Ramirez, Pitchayapa Jingjit, Becky Parker
Location:
12 N 519235 4251489
Time: departure at 10:00 am. Estimated return at 12:00 am
Duration: 2 hours approximately
Vehicles: 2 ATV’s and 1 Rover
Purposes:
Geology EVA.  Teach students about sedimentary rocks, deposition, lithification, and examination of the regolith.  This is the second half of EVA #4 that was shortened due to the NPR reporter visit.
Goal: Complete the Geology EVA.
Summary:
The location for this EVA was chosen in order to study different kinds of geologic structures and the methods by which they were created. The pattern of lithification in the sedimentary rock field was unique and interesting to the crew. Former evidence of a vigorous fluvial environment was observed at the second location. Determination of width and height were performed during the EVA.
Otsmar Villarroel
Crew 177 Commander

Journalist Report – March 30th

Journalist Report by Victoria LaBarre
March 30th/2017
This morning our crew was ecstatic to discover that Shannon had gone ahead and brought a Port-a-Potty for us to use the pipes underneath the Hab had been completely disconnected by a previous crew, rendering our sewage system completely useless. (We have worked out a system of discarding waste in bags and dumping them outside in the trash cans by the engineering airlock before Shannon brought us the Port-a-Potty. Exaggeration or not, there has never been a group of people so excited to receive a Port-a-Potty in the history of all mankind.
Soon after the crew broke in the Port-a-Potty, the first half of the crew went out for our first EVA. This EVA is the second half of the Geology EVA that had been cut short due to the NPR journalist’s presence and will also give our geologist a second chance to measure the height of the Little Canyon.
The two crew members working on the robot took the gearboxes off the wheels and adjusted the robot to prepare for our afternoon EVA where our crew will test the strength and dexterity of the robotic arm by picking up different sized rocks out near Cow Patty field. The robot was only able to pick up one rock due to damage from riding in the back or the rover and will need to be fixed before testing it again. The rock picked up by the robotic arm was taken back to the Hab to be weighed in the lab afterward.
The half of the crew out on EVA was excited to return to a working clock in the airlock. Our geologist/health and safety officer’s, Caleb Li, project to make a fully functioning clock was successful, and he installed the clock on the side of the front door airlock for the EVA crew members to use.
While one-half of the crew was out on EVA, our microbiologist began and finished the gram stain process for the bacterial samples taken previously and attempted to use the microscope.
After returning to the Hab, the crew gathered around our greenhouse officer’s bike generator project and watched as one of our crew members rode it so fast that he reached the 50 calories burned mark in 3 minutes.
After dinner, the crew spent a few hours helping coming up with a marketing plan to promote MDRS at our community college, McLennan Community College.

Commander Report – March 30th

Sol 5/March 30th, 2017
Crew Commander Report by Otsmar Villarroel, Crew 177 Commander
Although today was not as beautiful as previous days with a beautiful blue sky and calm wind, it was a great day for the crew’s mission here on the red planet. After breakfast, a group consisting of Parker, Li, Espinoza, and Jingjit left the Hab to conduct some geology experiments.
The main goal of this EVA was to study the terrain from a geological point of view and thus collect some geology data of interest.  The location for this EVA was chosen in order to study different kinds of geologic structures and the methods by which they were created. The pattern of lithification in the sedimentary rock field was unique and interesting to the crew.
Former evidence of a vigorous fluvial environment was observed at the second location. Determination of width and height were performed during the EVA. After collection of the data, the Geology EVA group returned to the Hab.
After enjoying lunch, at about 2:30 pm, Martian time, Crew members LaBarre and Espinoza, had the opportunity of testing her Emergency Medical Service Rover (EMRS) outside the Hab. The EMSR is a robot designed and programmed to retrieve an injured astronaut out on an EVA at Mars and bring him or her back to the Hab for medical assistance. During the test LaBarre and Espinoza detected some operational/mechanical issues.  However, they plans correcting them and a new test, pending approval, will be performed tomorrow.
Engineer, Espinoza, completed their daily routine without anomalies to be reported.  Greenhab officer, Ramirez, worked on taking care of plants as well as on his individual project, the bike generator.  He is planning to collect final results tomorrow to be analyzed back in Waco, Texas.  Our microbiologist, Jingjit, began and finished the gram stain process for the bacterial samples taken previously and attempted to use the microscope. Crew member, Li, spent part of the afternoon working on his clock.
After dinner, crew spent the crew spent a few hours helping coming up with a marketing plan to promote MDRS at our community college, McLennan Community College.  Tomorrow, pending approval, an EVA is planned. The purpose of this EVA is the collection of soil samples surrounding the Mars Desert Research Station (MDRS) to bring back to McLennan Community College in Waco, Texas in order to find bacteria producing antibiotics to combat antibiotic resistance crisis.
Last night, our crew astronomer, Becky Parker, used the telescope, with the main goal of completing a two-star alignment as well as a four- star calibration procedure. Unfortunately, while she was able to have the telescope working properly, she could not get the telescope to focus as expected.  We are planning to contact Mission Support looking for recommendations.

Astronomy Report – March 30th

Crew 177 Astronomer Report 30Mar2017
Astronomer Report
On Sol 1 & 2 the sky was too cloudy to attempt to use the telescope.  On Sol 3, we enjoyed the night sky with the naked eye.  On Sol 4 we went to the Musk Observatory and followed the quick guide procedures.
We were able to view and focus on Jupiter, but were unable to focus on Polaris or Beetle Juice.  Because we were unable to complete the alignment procedure we discontinued use of the telescope.
Becky Parker
Crew 177 Astronomer

Sol Summary – March 29th

Crew 177 Sol 4 Summary 29Mar2017
Person filling out Report:
Pitchayapa Jingjit—a biologist and a photographer for Crew 177
Summary Title:
Sol 4 at Mars
Mission Status:
Everyone is doing great and is hopeful about the toilet situation
SOL Activity Summary:
Improving individual/group research project and continuing the chemistry EVA. Depending on tonight’s sky, astronomy EVA may be conducted.
Look ahead Plan:
Continuing the geology EVA at the Little Canyon and depending on tonight’s sky, the astronomy EVA may be performed, and progressing through with individual/group projects.
Anomalies in Progress:
N/A
Weather:
The weather is sunny in the morning and afternoon. We are expecting a clear sky tonight.
Crew Physical Status:
Everyone is better adjusted at Mars and is happy that the plumber came to the HAB.

 

EVA Report – March 29th

EVA Report:  EVA #6
Crew members: Elijah Espinoza, Caleb Li, Esteban Espinoza, Becky Parker
Location: 12 N 520446 4251115 (Candor Chasma)
Time: departure at 9:00 am. Estimated return at 11:30 am
Duration: 2 hours and 30 minutes approximately
Vehicles: 2 ATV’s and 1 Rover
Purposes:
Chemistry EVA.  Students will be visiting the location known as Candor Chasma with the goal of collecting samples as well collect water sample.  All samples will be analyzed in the Science Lab paying close attention to the chemical composition and Teach students about using the Geiger counter to identify background radiation.  Also have students collect a water sample for analysis back at the hab.
Goal: Complete the Chemistry EVA.
Summary:
Unable to use the Geiger counter due undetermined technical issues. No water was found. For fauna, we observed birds, and found what appeared to be several nests, probably by a small mammal. Flora observed were flowers, grass, trees, cacti and brush. There were also many varieties of gypsum.  Some rock samples were brought back to the Hab for future chemical composition analysis.
Otsmar Villarroel
Crew 177 Commander